A spoon rides a roller coaster at the fair. He and his car have a mass of 500.0 kg and coast down a 100 m hill from rest. At the bottom of the hill he (and her car) collide with a car and riders, a ghost and a calf (m=850kg), that sit at rest. If the collision is elastic, will the ghost and the calf make it over the hill that is 50.0 m high?


1
Expert's answer
2021-03-08T08:28:48-0500
"mv=-mv'+Mu\\m\sqrt{2gH}=-mv'+Mu"

"u=\frac{2m}{m+M}\sqrt{2gH}\\
u=\frac{2(500)}{500+850}\sqrt{2(9.8)(100)}=32.8\frac{m}{s}"

And the needed speed will be


"U=\sqrt{2gh}=\sqrt{2(9.8)(50)}=31.3\frac{m}{s}<u"

Thus, the ghost and the calf will make it over the hill that is 50.0 m high.